if the slope of a line containing the points (3,a) and (b,3) is 2, what is a in terms of b?
Answer by stanbon(75887)

if the slope of a line containing the points (3,a) and (b,3) is 2, what is a in terms of b?
----
slope: (a-3)/(3-b) = 2
a-3 = 6-2b
a = -2b+9
==============
Cheers,
stan H.